OpenStudy (anonymous):

Use 3 and 4 as the first two terms of a Fibonacci-type sequence. (a) Find the next five terms? 3, 4, ___7____, ___11____, ____18___, ___29____, ____47___ (b) Write a symbolic description for this Fibonacci-type sequence using variables tn-1 and tn+2 as terms of the sequence.

OpenStudy (anonymous):

are 7, 11, 18, 29, 47 correct

OpenStudy (anonymous):

Those are the correct for part a for part b though, did you mean \[t _{n-1} \] & \[t _{n-2}\]?

OpenStudy (anonymous):

That is what the teacher wrote I have not answered that part yet, it that incorrect?

OpenStudy (anonymous):

if the teacher did want t_n in terms of t_(n-1) & t_(n+2), then the problem becomes a bit tricky. The easiest way to solve analytically is to use the fibonacci sequence definition of\[t _{n}=t _{n-1}+t _{n-2}\] for n+2 and n+1:\[t_{n+2}=t_{n+1}+t_n\]\[t_{n+1}=t_n+t_{n-1}\] and solve both of those in terms of t_(n+1)\[t_{n+1}=t_{n+2}-t_n\]\[t_{n+1}=t_n+t_{n-1}\]. Setting those equal to each other and solving for t_n we get\[t_{n+2}-t_n=t_n+t_{n-1}\]\[t_{n+2}-t_{n-1}=2t_n\]\[t_n=(t_{n+2}-t_{n-1})/2\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

which part are you confused about?

OpenStudy (anonymous):

I confused myself because I was looking at both formulas thinking it was solved instead of realzing they are two seperate terms

OpenStudy (anonymous):

Do you understand it now, or do i need to clarify a step or two?

OpenStudy (anonymous):

How did you get the n+2 and n+1?

OpenStudy (anonymous):

n+1 and n+2 are representing the next two terms, respectively. as far as how i derived the two formulae when i defined t_{n+1} and t_(n+2}, by taking the original Fibonacci function: \[t_n=t_{n-1}+t_{n-2}\] and shifting it over two. step 1: replace n with another letter, I will use p. \[t_p=t_{p-1}+t_{p-2}\] step 2: shift two to the + by setting n+2=p \[t_{n+2}=t_{n+2-1}+t_{n+2-2}=t_{n+1}+t_n\] you can use the same process for shifting it 1 to get the second equation i listed. hope that helps (:

OpenStudy (anonymous):

Yes this help a lot thank you